# Higher Gaussian maps on the hyperelliptic locus and second fundamental form

### v2

@@ 1. Introduction
where $H^{0}(C,K_{C}^{\otimes 2})^{+}$ denotes the invariant subspace of $H^{0}(C,K_{C}^{\otimes 2})$ under the action of the hyperelliptic involution. The canonical map is the composition of the $2:1$ map $\pi:C\rightarrow{\mathbb{P}}^{1}$ with the $(g-1)-{th}$ Veronese embedding $\nu_{g-1}:{\mathbb{P}}^{1}\rightarrow{\mathbb{P}}^{g-1}$, hence the space $I_{2}$ is identified with the vector space of quadrics containing the rational normal curve $\nu_{g-1}({\mathbb{P}}^{1})$ in ${\mathbb{P}}^{g-1}$. The second Gaussian map $\mu_{2}$ is a linear map
@@
If $C$ is a hyperelliptic curve, it was proven in [17], [3] that the rank of $\mu_{1}$ is $32g-$ 3, while in [4] it was shown that the rank of $\mu_{2}$ is $52g-$ 5.
@@
(See Theorem 3.1) Let $C$ be a hyperelliptic curve of genus $g\geq 3$ Then for every $0\leq k\leq\frac{g-1}{2}$
@@
Then, $Rank(\mu_{2k})=0$ for every $k>\lfloor\frac{g-1}{2}\rfloor$.
@@
where $N^{*}_{\mathcal{H}_{g}|\mathcal{A}_{g}}$ denotes the conormal bundle of $j_{h}(\mathcal{H}_{g})$ in $\mathcal{A}_{g}$. At a point $[C]\in{\mathcal{H}}_{g}$, $\rho_{HE}$ is therefore a linear map
@@
In [5] it was proven that the map $\rho_{HE}$ can be expressed in terms of the Hodge Gaussian map $\rho$ introduced in [8]. An important property proven in [8] is a formula computing, for any $Q\in I_{2}$, the Hodge Gaussian map $\rho(Q)$ on Schiffer variations at points $p$ on any curve $C$ in terms of the second Gaussian map $\mu_{2}(Q)$.
@@
In the case of hyperelliptic curves, the result of [8] gives the following expression for $\rho_{HE}(Q)$ on the Schiffer variation $\xi_{p}\in H^{1}(C,T_{C})^{+}$ at a Weierstrass point $p\in C$, in terms of the second Gaussian map :
$\rho_{HE}(Q)(\xi_{p}\odot\xi_{p})=-2\pi i\mu_{2}(Q)(p).$
Then we have $\rho_{HE}(Q)(\xi_{p}\odot\xi_{p})=0$, since $p$ is a Weierstrass point and one can easily see that for all the quadrics $Q\in I_{2}$, we have $\mu_{2}(Q)(p)=0$ (see Remark 6.13).
@@
Here we exploit similar techniques to compute $\rho_{HE}(Q)$ for quadrics $Q$ contained in the kernel of higher Gaussian maps on odd higher Schiffer variations $\xi_{p}^{2k+1}\in H^{1}(C,T_{C})^{+}$ at a Weierstrass point $p\in C$ (see Section 4 for the definition of higher Schiffer variations). For any $k\leq 2g-3$, consider the $(k+1)$-dimensional subspace
@@
Clearly tangent directions to totally geodesic subvarieties are asymptotic directions. Notice that saying that a nonzero element $\zeta\in H^{1}(T_{C})^{+}$ is an asymptotic direction means that the point $[\zeta]\in{\mathbb{P}}H^{1}(T_{C})^{+}$ is in the base locus of the linear space of quadrics $\rho_{HE}(I_{2})\subset Sym^{2}(H^{1}(T_{C})^{+})^{\vee}$. In [5] it is shown that $\rho_{HE}$ is injective, hence the image of $\rho_{HE}$ in $Sym^{2}(H^{1}(T_{C})^{+})^{\vee}$ has dimension $\frac{(g-1)(g-2)}{2}$. So $\rho_{HE}(I_{2})$ is a linear system of quadrics in ${\mathbb{P}}(H^{1}(C,T_{C})^{+})\cong{\mathbb{P}}^{2g-2}$ of dimension $\frac{(g-1)(g-2)}{2}$.
@@
As a consequence we have the following
Remark 1.4.
(see Remark 6.15).
In [5, Corollary 3.4] it is shown that $\rho_{HE}$ is injective, so
for all $k$. For any hyperelliptic curve $C$ of genus $g\geq 3$ and for any Weiestrass point $p\in C$, we have proven that we have the following chain of subspaces of quadrics in ${\mathbb{P}}H^{1}(T_{C})^{+}$ and of corresponding maximal isotropic subspaces in $V_{\frac{g-3}{2}}$ if $g$ is odd and in $V_{\frac{g-2}{2}}$ if $g$ is even.
If $g$ is odd:
while if $g$ is even we have
@@ 2. Gaussian maps
where $I_{2}(L)$ is the kernel of the multiplication map $S^{2}H^{0}(C,L)\rightarrow H^{0}(C,L^{\otimes 2})$. Since $\Phi_{1,L}$ vanishes on symmetric tensors, one writes
@@
If $\alpha=\sum l_{i}\otimes m_{i}\in I_{2}(L)$ with $l_{i}=f_{i}(z)l$ and $m_{i}=g_{i}(z)l$, using that $\mu_{0,L}(\alpha)=0$, and $\Phi_{1,L}(\alpha)=0$, one finds the identities
(4) $\sum f_{i}g_{i}\equiv 0,\ \ \sum f_{i}^{\prime}g_{i}-f_{i}g_{i}^{\prime}\equiv 0.$
Taking derivatives of the first equation in (4) and using the second one we obtain the identities:
(5) $\sum f_{i}^{\prime}g_{i}\equiv\sum f_{i}g_{i}^{\prime}\equiv 0,$
Derivating the equations in (5) we have:
(6) $\sum f_{i}^{\prime\prime}g_{i}=-\sum f_{i}^{\prime}g_{i}^{\prime}=\sum f_{i}g_{i}^{\prime\prime}.$
The local expression of $\mu_{2,L}$ is
(7) $\mu_{2,L}(\alpha)=\sum f_{i}^{\prime\prime}g_{i}\ dz^{\otimes 2}\otimes l^{\otimes 2}=\sum f_{i}g_{i}^{\prime\prime}\ dz^{\otimes 2}\otimes l^{\otimes 2}=-\sum f_{i}^{\prime}g_{i}^{\prime}dz^{\otimes 2}\otimes l^{\otimes 2}.$
More generally the local expression for $\Phi_{2k,L}$ when $k\geq 1$ is given by
(8) $\Phi_{2k,L}(\alpha)={(-1)^{k}}\sum f_{i}^{(k)}g_{i}^{(k)}\ dz^{\otimes 2k}\otimes l^{\otimes 2},$
and since it vanishes on antisymmetric tensors one defines
(9) $\mu_{2k,L}:=\Phi_{2k,L_{|_{Ker(\mu_{2k-2,L})}}}.$
Analogously the local expression for $\Phi_{2k+1,L}$ when $k\geq 0$ is given by
(10) $\Phi_{2k+1,L}(\alpha)=\sum(f_{i}^{(k+1)}g_{i}^{(k)}-f_{i}^{(k)}g_{i}^{(k+1)})\ dz^{\otimes 2k+1}\otimes l^{\otimes 2},$
and since it vanishes on symmetric tensors one defines for every $k\geq 2$
(11) $\mu_{2k+1,L}:=\Phi_{2k+1,L_{|_{Ker(\mu_{2k-1,L})}}}.$
Moreover since $\alpha=\sum l_{i}\otimes m_{i}\in Ker(\mu_{2k-2,L})$ if and only if
(12) $\mu_{0,L}(\alpha)=\Phi_{0,L}(\alpha)=...=\mu_{2k-2,L}(\alpha)=\Phi_{2k-1,L}(\alpha)=0,$
and these conditions give identities
(13) $\sum f_{i}^{(h)}g_{i}^{(r)}\equiv 0,\forall h,r\ \text{such that}\ h+r\leq 2k-1,$
the local expression for $\mu_{2k,L}$ is equivalent to
(14) $\mu_{2k,L}(\alpha)=(-1)^{m}\sum f_{i}^{(2k-m)}g_{i}^{(m)}dz^{\otimes 2k}\otimes l^{\otimes 2},$
for every $m=0,...,2k$. Analogously $\alpha=\sum l_{i}\otimes m_{i}\in Ker(\mu_{2k-1,L})$ if and only if
(15) $\Phi_{0,L}(\alpha)=\mu_{1,L}(\alpha)=...=\mu_{2k-1,L}(\alpha)=\Phi_{2k,L}(\alpha)=0,$
and these conditions give the identities
(16) $\sum f_{i}^{(h)}g_{i}^{(r)}\equiv 0,\forall h,r\ \text{such that}\ h+r\leq 2k,$
and from this, it follows that we can equivalently express (10) as
(17) $\mu_{2k+1,L}(\alpha)=(-1)^{m}\sum(f_{i}^{(2k+1-m)}g_{i}^{(m)}-f_{i}^{(m)}g_{i}^{(2k+1-m)})dz^{\otimes 2k+1}\otimes l^{\otimes 2}$
for every $m=0,...,k$.
@@ 3. Rank of higher order Gaussian maps on hyperelliptic curves
and $M\cong L^{\otimes(g-2)}$. Then $H^{0}(C,M)\cong H^{0}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}(g-2))$ has dimension $g-1$.
@@
Let $C$ be a hyperelliptic curve of genus $g\geq 3$ Then for every $0\leq k\leq\frac{g-1}{2}$
@@
Now take $n\geq 2$ and suppose that $(ii)$ holds for every $10\leq k<n-$ 1 and $(i)$ holds for every $1\leq k<n$. We are going to prove that $(ii)$ holds for $k=n-1$, which automatically implies that $(i)$ holds for $k=n$. Set $k=n-1$ and suppose that
@@
$\omega_{i}=f_{i}dz^{\prime},\ t=fT\ \text{and}\ s=gT,$
where $f_{i}$ is a holomorphic function, $dz^{\prime}$ is a local generator on $M=K_{C}\otimes L^{\vee}$ and $T$ is a local generator on $L$. From the definition of $Q_{ij}$ in Lemma 3.3, it follows that we can write locally $Q$ as
@@
$-\left[\sum\limits_{1\leq i<j\leq g-1}a_{ij}\sum\limits_{r=0}^{k+1}\binom{k+1}{h}f^{(k+1-h)}f_{j}^{(h)}\sum\limits_{e=0}^{k+1}\binom{k+1}{l}g^{(k+1-l)}f_{i}^{(l)}\right]dz^{\otimes(2k+4)}$
@@
$Q_{ij}$ as in 3.3, with respect to the basis of $H^{0}(M)$ given by $\alpha_{i}=x^{i}\frac{dx}{y}$, for $i=1,...,g-1$. Then $Q\in Ker(\mu_{2})$ if and only if for all $33\leq l\leq 2g-$ 3
@@
with $Q_{ij}=s\alpha_{i}\odot t\alpha_{j}-s\alpha_{j}\odot t\alpha_{i}$ as in 3.3, for $i=1,...,g-1.$ For any $k\geq 2$, $Q\in Ker(\mu_{2k})$ if and only if
@@
We will argue locally. Take the basis of $H^{0}(C,K_{C})$ given in (30) and the one for $H^{0}(C,M)$ given in (31). We will proceed by induction on $k$. The thesis holds for $k=1$ by Lemma 3.6. Now we suppose that the thesis holds for $k-1$, $k\geq 2$ and take $Q\in Ker(\mu_{2k-2})$. Since $Q$ also belongs to the previous kernels, proving the lemma is equivalent to prove that $Q\in Ker(\mu_{2k})$ if and only if
@@
for all $32k-1\leq l\leq 2g-$ 3. Using lemma 3.5 we have that $\mu_{2k}(Q)=0$ if and only if $\mu_{2k-1,L}(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}\wedge\omega_{j}))=0$, that is if and only if
@@
So we have shown that for $Q\in Ker\mu_{2k-2}$, $\mu_{2k}(Q)=0$ if and only if for every $32k-1\leq l\leq 2g-$ 3
@@
We will start introducing some notations. For every $i,j$ such that $11\leq i<j\leq g-$ 1, for every $l=3,...,2g-3$, let $(w_{l})^{2}$ the vector whose coordinates are
@@
which are the coefficients of the $l$-th equation of the set of equations given in Lemma 3.7, describing $Ker(\mu_{2})$, ordered by increasing values of $i$. For every $2\leq r\leq k$ and $l=2r-1,...,2g-3$, let $w_{l}^{(2r)}$ be the vector whose coordinates are
@@
Observe that $B^{\prime}_{k,l}$ is just the matrix with rows $w_{l}^{(2)},...,w_{l}^{(2k)}$ where we have removed the entries corresponding to $i+j\neq l$. Notice that these entries are all 00 for all the vectors $w_{l}^{(2r)}$.
Define $c_{k,l}=min\{n_{l},k\}$ and let $B_{k,l}$ be the minor of $B^{\prime}_{k,l}$ of order $c_{k,l}$ given by the first $c_{k,l}$ rows and $c_{k,l}$ columns. We are going to prove that it is not zero for all $33\leq l\leq 2g-$ 3 by induction on $k$.
Consider first the case $k=2$. Observe that if $l=3,4,2g-4,2g-3$, then $n_{l}=1$. In this case $B_{2,l}=j-i>0$ where $(i,j)$ is the only pair such that $i+j=l$. If $55\leq l\leq 2g-$ 5 then $n_{l}\geq 2$ and $c_{2,l}=min(2,n_{l})=2$ and we have
@@
whose determinant is $j-i-1>0$. Now assume that for every $2\leq k_{0}\leq k$ and for every $33\leq l\leq 2g-$ 3, $B_{k,l}$ is not zero. We want to prove that the same holds for $k+1$. If $n_{l}\leq k$, then $n_{l}<k+1$, $c_{k+1,l}=min\{n_{l},k\}=n_{l}$ and $B_{k+1,l}=B_{k,l}$. In this case, we have nothing to prove. Assume then $n_{l}>k$. Then $c_{k+1,l}=min\{n_{l},k+1\}=k+1$. Then $B_{k+1,l}$ is the following $(k+1)\times(k+1)$ matrix
$\begin{pmatrix}j-i&j-i-2&...&j-i-2k\\ (j-i)ji&(j-2-i)(i+1)(j-1)&...&(j-i-2k)(j-k)(i+k)\\ .&.&.&.\\ .&.&.&.\\ .&.&.&.\\ (j-i)ji...(j-(k+1-2))(i-(k+1-2))&&...&(j-i-2k)(j-k)(i+k)...(j-2k+1)(i+1)\end{pmatrix}$
where $i$ is the minimum such that $i+j=l$ and where we mean that every entry in the $r$-th row, $11\leq r\leq k+$ 1 (corresponding to an index $(i,j)$) is zero whenever $i\leq r-1$.
The determinant of the above matrix is equal to the product of $(j-i)...(j-i-2k)$ by the determinant of the following matrix
$\begin{pmatrix}1&1&...&1\\ ji&(i+1)(j-1)&...&(j-k)(i+k)\\ .&.&.&.\\ .&.&.&.\\ .&.&.&.\\ (j-i)ji...(j-(k+1-2))(i-(k+1-2))&&...&(j-k)(i+k)...(j-2k+1)(i+1)\end{pmatrix}$
@@
$\begin{pmatrix}(j-1)-i&...&(j-1)-i-2(k-1)\\ .&.&.\\ .&.&.\\ .&.&.\\ (j-1-i)(j-1)i...((j-1-(k-2)))(i-((k-2)))&...&(j-1-i-2(k-1))(j-1-(k-1))(i+k-1)...(j-2(k-1)+1)(i+1)\end{pmatrix}.$
is $B_{k,l-1}$ (where as we have already said, we mean that every entry corresponding to a index $(i,j)$ in the $r$-th row, $1\leq r\leq k$ is zero whenever $i\leq r-1$. The matrix has not zero determinant by the inductive hypothesis. Hence we have shown, by induction, that for any $33\leq l\leq 2g-$ 3 and for any $k\geq 2$, $(w_{l})^{2}$,…,$(w_{l})^{2k}$ impose exactly $c_{k,l}$ linearly independent conditions.

### v3

@@ 1. Introduction
where $H^{0}(C,K_{C}^{\otimes 2})^{+}$ denotes the invariant subspace of $H^{0}(C,K_{C}^{\otimes 2})$ under the action of the hyperelliptic involution. The canonical map is the composition of the $2:1$ map $\pi:C\rightarrow{\mathbb{P}}^{1}$ with the (g1)th(g-1)-{th} Veronese embedding $\nu_{g-1}:{\mathbb{P}}^{1}\rightarrow{\mathbb{P}}^{g-1}$, hence the space $I_{2}$ is identified with the vector space of quadrics containing the rational normal curve $\nu_{g-1}({\mathbb{P}}^{1})$ in ${\mathbb{P}}^{g-1}$. The second Gaussian map $\mu_{2}$ is a linear map
@@
If $C$ is a hyperelliptic curve, it was proven in [17], [3] that the rank of $\mu_{1}$ is 2g32g-3, while in [4] it was shown that the rank of $\mu_{2}$ is 2g52g-5.
@@
(See Theorem 3.1) Let $C$ be a hyperelliptic curve of genus $g\geq 3$ . Then for every $0\leq k\leq\frac{g-1}{2}$
@@
Then, for every $k>\lfloor\frac{g-1}{2}\rfloor$ the domain of $\mu_{2k}$ is 0, hence $Rank(\mu_{2k})=0$.
@@
where $N^{*}_{\mathcal{H}_{g}|\mathcal{A}_{g}}$ denotes the conormal bundle of $j_{h}(\mathcal{H}_{g})$ in $\mathcal{A}_{g}$. At a point $C]\in{\mathcal{H}}_{g}$, $\rho_{HE}$ is therefore a linear map
@@
In [5] it was proven that the map $\rho_{HE}$ can be expressed in terms of the Hodge Gaussian map $\rho$ introduced in [8] and that it is injective. An important property proven in [8] is a formula computing, for any $Q\in I_{2}$, the Hodge Gaussian map $\rho(Q)$ on Schiffer variations at points $p$ on any curve $C$ in terms of the second Gaussian map $\mu_{2}(Q)$.
@@
In the case of hyperelliptic curves, the result of [5, section 5] gives an expression for $\rho_{HE}(Q)$ on the Schiffer variation $\xi_{p}\in H^{1}(C,T_{C})^{+}$ at a Weierstrass point $p\in C$, in terms of the second Gaussian map. This allows to show that $\rho_{HE}(Q)(\xi_{p}\odot\xi_{p})=0$, for any Weierstrass poin $p$ (see Remark 6.13).
@@
Here we exploit similar techniques to compute $\rho_{HE}(Q)$ for quadrics $Q$ contained in the kernel of higher Gaussian maps on certain tangent directions given by odd higher Schiffer variations $\xi_{p}^{2k+1}\in H^{1}(C,T_{C})^{+}$ at a Weierstrass point $p\in C$. We refer to Section 4 for the definition of higher Schiffer variations. In a local coordinate centered at a point $p$, the element $\xi_{p}^{n}\in H^{1}(C,T_{C})$ has a Dolbeault representative given by $\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}$, where $\rho_{p}$ is a bump function in $p$ which is equal to one in a small neighborhood $U$ containing $p$. If $p$ is a Weierstrass point and $n$ is odd, one immediately sees that $\xi_{p}^{n}$ is invariant under the hyperelliptic involution, hence it is tangent to the hyperelliptic locus.
So, for any $k\leq 2g-3$, consider the $(k+1)$-dimensional subspace
@@
Clearly tangent directions to totally geodesic subvarieties are asymptotic directions. Notice that saying that a nonzero element $\zeta\in H^{1}(T_{C})^{+}$ is an asymptotic direction means that the point $\zeta]\in{\mathbb{P}}H^{1}(T_{C})^{+}$ is in the base locus of the linear space of quadrics $\rho_{HE}(I_{2})\subset Sym^{2}(H^{1}(T_{C})^{+})^{\vee}$. In [5] it is shown that $\rho_{HE}$ is injective, hence the image of $\rho_{HE}$ in $Sym^{2}(H^{1}(T_{C})^{+})^{\vee}$ has dimension $\frac{(g-1)(g-2)}{2}$. So $\rho_{HE}(I_{2})$ is a linear system of quadrics in ${\mathbb{P}}(H^{1}(C,T_{C})^{+})\cong{\mathbb{P}}^{2g-2}$ of dimension $\frac{(g-1)(g-2)}{2}$.
@@ 2. Gaussian maps
where $I_{2}(L)$ is the kernel of the multiplication map $\mu_{0,L}:S^{2}H^{0}(C,L)\rightarrow H^{0}(C,L^{\otimes 2})$. Since $\Phi_{1,L}$ vanishes on symmetric tensors, one writes
@@
In general, since even Gaussian maps vanish on antisymmetric tensors and odd Gaussian maps vanish on symmetric tensors, for all $k\geq 1$, we set
(4) $\mu_{2k,L}:={\Phi_{2k,L}}_{|_{Ker(\mu_{2k-2,L})}}:Ker(\mu_{2k-2,L})\rightarrow H^{0}(L^{\otimes 2}\otimes K_{C}^{\otimes 2k}).$
(5) $\mu_{2k+1,L}:={\Phi_{2k+1,L}}_{|_{Ker(\mu_{2k-1,L})}}:Ker(\mu_{2k-1,L})\rightarrow H^{0}(L^{\otimes 2}\otimes K_{C}^{\otimes 2k+1}).$
We will now give a local expression for these Gaussian maps.
Choose a basis $\eta_{1},...,\eta_{r}$ of $H^{0}(L)$, whose local expressions is $\eta_{i}=f_{i}(z)l$, where $l$ is a local frame for $L$. Assume now that $Q=\sum_{ij}a_{ij}\eta_{i}\otimes\eta_{j}\in I_{2}(L)$, with $a_{ij}=a_{ji}$. Then we have:
$\sum_{ij}a_{ij}f_{i}(z)f_{j}(z)=0,$
since $Q\in I_{2}(L)$. Derivating we get $\sum_{ij}a_{ij}f^{\prime}_{i}(z)f_{j}(z)=0,$ and taking another derivative, we obtain
$\sum_{ij}a_{ij}f^{(2)}_{i}(z)f_{j}(z)=-\sum_{ij}a_{ij}f^{\prime}_{i}(z)f^{\prime}_{j}(z).$
Then
$\mu_{2,L}(Q)=\sum_{ij}a_{ij}f^{(2)}_{i}(z)f_{j}(z)dz^{\otimes 2}\otimes l^{\otimes 2}=-\sum_{ij}a_{ij}f^{\prime}_{i}(z)f^{\prime}_{j}(z)dz^{\otimes 2}\otimes l^{\otimes 2}.$
In general $Q=\sum_{ij}a_{ij}\eta_{i}\otimes\eta_{j}\in Ker(\mu_{2k-2,L})$ if and only if
(6) $\sum_{i,j}a_{ij}f_{i}^{(h)}(z)f_{j}^{(n)}(z)\equiv 0,\ \ \forall h,n,\ \text{such that }\ h+n\leq 2k-1.$
Moreover, $\forall n=0,...,2k$, we have:
(7) $\mu_{2k,L}(Q)=(-1)^{n}\sum_{i,j}a_{ij}f_{i}^{(2k-n)}(z)f_{j}^{(n)}(z)dz^{\otimes 2k}\otimes l^{\otimes 2}.$
Analogously, $\alpha=\sum_{i<j}a_{ij}(\eta_{i}\wedge\eta_{j})\in Ker(\mu_{2k-1,L})$ if and only if
$\sum_{i<j}a_{ij}(f_{i}^{(h)}(z)f_{j}^{(n)}-f_{i}^{(n)}(z)f_{j}^{(h)})(z)\equiv 0,\ \ \forall h,n,\ \text{such that }\ h+n\leq 2k$
and we have
(8) $\mu_{2k+1,L}(\alpha)=(-1)^{n}\sum_{i<j}a_{ij}(f_{i}^{(2k+1-n)}(z)f_{j}^{(n)}(z)-f_{j}^{(2k+1-n)}(z)f_{i}^{(n)}(z))dz^{\otimes 2k+1}\otimes l^{\otimes 2}.$
for every n=0,,2k+1n=0,...,2k+1.
@@ 3. Rank of higher order Gaussian maps on hyperelliptic curves
and $M\cong L^{\otimes(g-2)}$. Then $H^{0}(C,M)\cong H^{0}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}(g-2))$ has dimension g1g-1.
@@
Let $C$ be a hyperelliptic curve of genus $g\geq 3$ . Then for every $0\leq k\leq\frac{g-1}{2}$
@@
Now take $n\geq 2$ and suppose that $(ii)$ holds for every $0\leq k<n-1$ and $(i)$ holds for every $1\leq k<n$. We are going to prove that $(ii)$ holds for k=n1k=n-1, which automatically implies that $(i)$ holds for $k=n$. Set k=n1k=n-1 and suppose that
@@
$\omega_{i}=f_{i}dz\otimes T^{\vee},\ t=fT\ \text{and}\ s=gT,$
where $f_{i}$ is a holomorphic function and $T$ is a local generator on $L$. From the definition of $Q_{ij}$ in Lemma 3.3, it follows that we can write locally $Q$ as
@@
$-\left[\sum\limits_{1\leq i<j\leq g-1}a_{ij}\sum\limits_{h=0}^{k+1}\binom{k+1}{h}f^{(k+1-h)}f_{j}^{(h)}\sum\limits_{l=0}^{k+1}\binom{k+1}{l}g^{(k+1-l)}f_{i}^{(l)}\right]dz^{\otimes(2k+4)}$
@@
Set $S=C\times C$, and let $\Delta$ be the diagonal as usual. Then we have the following commutative diagram of exact sequences
[…]
So, taking global sections we get
(24) $H^{0}(S,(M\boxtimes M)(-k\Delta))$$H^{0}(S,(M\boxtimes M)(-k\Delta)_{|\Delta}$$H^{0}(S,(K_{C}\boxtimes K_{C})(-k\Delta))$$H^{0}(S,(K_{C}\boxtimes K_{C})(-k\Delta)_{|\Delta})$$\scriptstyle{\Phi_{k,M}}$$\scriptstyle{\Phi_{k}}$
Hence the corresponding diagram for the even order Gaussian maps:
(25) $Ker(\mu_{2k-2,M})$$H^{0}(C,M^{\otimes 2}\otimes K_{C}^{\otimes 2k})$$Ker(\mu_{2k-2})$$H^{0}(C,K_{C}^{\otimes 2k+2})$$\scriptstyle{\mu_{2k,M}}$$\scriptstyle{\mu_{2k}}$
Hence to compute the Gaussian maps $\mu_{2k,M}$, we compute the restriction of the Gaussian maps of the canonical bundle $\mu_{2k}$ to the subspace $Ker(\mu_{2k-2,M})$ of $Ker(\mu_{2k-2})$.
@@
$Q_{ij}$ as in 3.3, with respect to the basis of $H^{0}(M)$ given by $\alpha_{i}=x^{i}\frac{dx}{y}$, for i=1,,g1i=1,...,g-1. Then $Q\in Ker(\mu_{2})$ if and only if for all $3\leq l\leq 2g-3$
@@
with $Q_{ij}=s\alpha_{i}\odot t\alpha_{j}-s\alpha_{j}\odot t\alpha_{i}$ as in 3.3, for i=1,,g1.i=1,...,g-1. For any $k\geq 2$, $Q\in Ker(\mu_{2k})$ if and only if
@@
We will argue locally. Take the basis of $H^{0}(C,K_{C})$ given in (21) and the one for $H^{0}(C,M)$ given in (22). We will proceed by induction on $k$. The thesis holds for $k=1$ by Lemma 3.6. Now we suppose that the thesis holds for k1k-1, $k\geq 2$ and take $Q\in Ker(\mu_{2k-2})$. Since $Q$ also belongs to the previous kernels, proving the lemma is equivalent to prove that $Q\in Ker(\mu_{2k})$ if and only if
@@
for all $2k-1\leq l\leq 2g-3$. Using lemma 3.5 we have that $\mu_{2k}(Q)=0$ if and only if $\mu_{2k-1,M}(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\alpha_{i}\wedge\alpha_{j}))=0$, that is if and only if
@@
So we have shown that for $Q\in Ker(\mu_{2k-2}$), $\mu_{2k}(Q)=0$ if and only if for every $2k-1\leq l\leq 2g-3$
@@
We will start introducing some notations. For every $i,j$ such that $1\leq i<j\leq g-1$, for every l=3,,2g3l=3,...,2g-3, let $(w_{l})^{2}$ the vector whose coordinates are
@@
which are the coefficients of the $l$-th equation of the set of equations given in Lemma 3.7, describing $Ker(\mu_{2})$, ordered by increasing values of $i$. For every $2\leq r\leq k$ and l=2r1,,2g3l=2r-1,...,2g-3, let $w_{l}^{(2r)}$ be the vector whose coordinates are
@@
Observe that $B^{\prime}_{k,l}$ is just the matrix with rows $w_{l}^{(2)},...,w_{l}^{(2k)}$ where we have removed the entries corresponding to $i+j\neq l$. Notice that these entries are all 0 for all the vectors $w_{l}^{(2r)}$.
Define $c_{k,l}=min\{n_{l},k\}$ and let $B_{k,l}$ be the minor of $B^{\prime}_{k,l}$ of order $c_{k,l}$ given by the first $c_{k,l}$ rows and $c_{k,l}$ columns. We are going to prove that it is not zero for all $3\leq l\leq 2g-3$ by induction on $k$.
Consider first the case $k=2$. Observe that if l=3,4,2g4,2g3l=3,4,2g-4,2g-3, then $n_{l}=1$. In this case $B_{2,l}=j-i>0$ where $(i,j)$ is the only pair such that $i+j=l$. If $5\leq l\leq 2g-5$ then $n_{l}\geq 2$ and $c_{2,l}=min(2,n_{l})=2$ and we have
@@
whose determinant is ji1>0j-i-1>0. Now assume that for every $2\leq k_{0}\leq k$ and for every $3\leq l\leq 2g-3$, $B_{k,l}$ is not zero. We want to prove that the same holds for $k+1$. If $n_{l}\leq k$, then $n_{l}<k+1$, $c_{k+1,l}=min\{n_{l},k\}=n_{l}$ and $B_{k+1,l}=B_{k,l}$. In this case, we have nothing to prove. Assume then $n_{l}>k$. Then $c_{k+1,l}=min\{n_{l},k+1\}=k+1$. Then $B_{k+1,l}$ is the following $(k+1)\times(k+1)$ matrix
(j -i j-i-2…j-i-2k (j -i)ji (j-2-i)(i+1)(j-1)…(j-i-2k)(j-k)(i+k) . . .. . . .. . . .. (j-i)ji…(j-(k+1-2))(i-(k+1-2)) …(j-i-2k)(j-k)(i+k)…(j-2k+1)(i+1) )
where $i$ is the minimum such that $i+j=l$ and where we mean that every entry in the $r$-th row, $1\leq r\leq k+1$ (corresponding to an index $(i,j)$) is zero whenever $i\leq r-1$.
The determinant of the above matrix is equal to the product of (ji)(ji2k)(j-i)...(j-i-2k) by the determinant of the following matrix
(1 1…1 ji (i+1)(j-1)…(j-k)(i+k) . . .. . . .. . . .. (j-i)ji…(j-(k+1-2))(i-(k+1-2)) …(j-k)(i+k)…(j-2k+1)(i+1) )
@@
((j-1)-i…(j-1)-i-2(k-1) . .. . .. . .. (j-1-i)(j-1)i…((j-1-(k-2)))(i-((k-2)))…(j-1-i-2(k-1))(j-1-(k-1))(i+k-1)…(j-2(k-1)+1)(i+1) ).
is $B_{k,l-1}$ (where as we have already said, we mean that every entry corresponding to a index $(i,j)$ in the $r$-th row, $1\leq r\leq k$ is zero whenever $i\leq r-1$. The matrix has not zero determinant by the inductive hypothesis. Hence we have shown, by induction, that for any $3\leq l\leq 2g-3$ and for any $k\geq 2$, $(w_{l})^{2}$$(w_{l})^{2k}$ impose exactly $c_{k,l}$ linearly independent conditions.